The resistance of a copper wire increases by a factor of 1.62 when current flows through the wire compared to its resistance when no current flows and its temperature is 20o C. What is its temperature when the current flows?

1
Expert's answer
2022-06-01T12:49:00-0400
"R(t_2)=R(t_1)(1+\alpha(t_2-t_1))"

"\frac{R(t_2)}{R(t_1)}=1+\alpha(t_2-t_1)"

"1.62=1+0.00386*(t_2-20)"

"t_2=181\:{^\circ\rm C}"